OpenStudy (anonymous):

Sigma Sum n=1,INFINITY[((n!)^2)/(kn!)] For which positive integers k is the series convergent?...I am stumped...

OpenStudy (agreene):

\[\sum_{n=1}^{\infty} \frac{n!^2}{kn!}\] I really dont remember my convergence tests... but it seems like the numerator will approach infinity much quicker than the denominator for all k not equal to 0. which would cause it to be perminately divergent except for k=0 where it is undefined.

OpenStudy (anonymous):

I thought k= infinity was right as well. it is wrong. The answer is something like k\[\le\]???

OpenStudy (anonymous):

Oops reverse that....

OpenStudy (anonymous):

i thought k = 0 ....not sure why it plopped infinity in there.

OpenStudy (anonymous):

Are you sure that the equation isn't \[\sum_{n=1}^{\infty} {{(n!)^{2}}\over{(kn)!}} \]

OpenStudy (anonymous):

Yeah, thats the equation

OpenStudy (anonymous):

Any ideas?

OpenStudy (anonymous):

Well it definitely converges for some k... Let me think about it for a bit

OpenStudy (anonymous):

thanks brah

OpenStudy (anonymous):

Well I'm pretty sure it converges for k >= 2

OpenStudy (anonymous):

sure does! THanks

OpenStudy (anonymous):

Whats the logic?

OpenStudy (anonymous):

So, n! grows a lot faster than n^2

OpenStudy (anonymous):

so (kn)! is applying the factorial after you multiply by a constant where (n!)^2 evaluates n! then applies the square

OpenStudy (anonymous):

The distance between consecutive factorial terms grows a lot faster than squaring terms

OpenStudy (anonymous):

But if k >= 2 then we are at least doubling that distance before we even apply the factorial

OpenStudy (anonymous):

10! << 20!

OpenStudy (anonymous):

(10!)^2 = (10!)(10!) < (20!) = (20*19*18*17*16*15*14*13*12*11)*10!

OpenStudy (anonymous):

hmmm well thanks! Gotta think about it more.

OpenStudy (jamesj):

Use the ratio test. The sum of an infinite series x_n converges if limit |x_{n+1} / x_n | < 1 as n --> infinity.